# Far from Perfect: Quantum Error Correction with (Hyperinvariant) Evenbly Codes

### v1

@@ E.1 Greedy Reconstruction Algorithm
A non-optimal but efficient decoder against erasure errors on holographic codes is given by a greedy algorithm, first proposed in [8] for HaPPY codes. In holographic codes, the logical qubits are arranged on a hyperbolic lattice, resulting in a hierarchy with respect to bulk depth, i.e., lattice distance from the physical qubits on the boundary. After some number of known (heralded) erasures of the physical qubits has occured, one starts to reconstruct the logical qubits closest to the boundary, applying the decoding map corresponding to each individual tensor (e.g., of the $\llbracket 5,1,3\rrbracket$ code for the perfect-tensor pentagon code). Once these are (partially) reconstructed, one has access to the “intermediate” physical qubits corresponding to contracted legs one layer deeper into the bulk, repeating the local reconstruction procedure there then repeating the procedure. When no more logical qubits can be reconstructed, the greedy algorithm terminates and one has effectively found a decoder for logical qubits in a greedy wedge of the bulk on the non-erased physical qubits. A greedy decoder can be straightforwardly constructed for any tensor network stabilizer code by translating possible operator pushing steps, which follow from the application of the stabilizers and gauge-fixed logical operators, into greedy steps. For any Evenbly code, for example, the operator pushing steps of Figure 5 (which are equivalent to the generalized isometry conditions of Figure 4) imply a greedy recovery for single and bulk-neighboring logical qubits that is shown in Figure 12(a)-(b). Note that this generalizes the original greedy algorithm of [8], in which each greedy step only recovers one logical qubit at a time. Beyond these general recovery step that any Evenbly code fulfills, the GHZ-based code discussed in the main text allows for an additional greedy step shown in Figure 12(c), whereby access to physical sites on a tensor can assist in the recovery of the logical qubit on its neighbor. Yet, greedy recovery for a max-rate Evenbly code is generally much weaker than for a HaPPY code, owing to the effect that low-weight errors can have deep in the bulk [24], precluding the HaPPY codes’ feature of “uberholography” [40]. A generic example is given in Figure 12(d), where the erasure of only a few boundary sites prevents greedy reconstruction of most of the bulk.
@@ E.2 Gaussian Elimination Algorithm
After obtaining the binary reduced-row echelon form (RREF) $BSV^{\prime}_{\text{RREF}}$, we iterate through the rows of $BSV^{\prime}_{\text{RREF}}$ and check the last two columns. If either of the entries are of value 1, then we check the other values in the other $(n-2)$ columns. If any value is equal to 1, then the algorithm succeeds, and we return $\mathrm{True}$ after incrementing the counter num_true. The algorithm terminates when all trials have completed, and the number of $\mathrm{True}$ values are counted against the number of trials (which we name $\mathcal{P}_{\text{rec}}=\frac{\text{num\_true}}{\text{mc\_trials}}$). This mathematical procedure largely follows [9].
@@
In the simulations executed for this work, we vary the number of Monte Carlo simulations as a function of the number of layers, in order to efficiently delegate computational resources. The number of Monte Carlo trials per layer and variant tested are shown in Table 2. The first column, labeled ”HTN”, corresponds to the zero-rate and constant-rate Evenbly codes tested in Sections III.5 and III.6, while the others refer to the other variants of the constant-rate HTN in which extra layers of Z or $Y$ gauge are added.
[…]
@@
[…]

### v2

@@ Appendix E Decoder Details
E.1 Greedy Reconstruction Algorithm
A non-optimal but efficient decoder against erasure errors on holographic codes is given by a greedy algorithm, first proposed in [8] for HaPPY codes. In holographic codes, the logical qubits are arranged on a hyperbolic lattice, resulting in a hierarchy with respect to bulk depth, i.e., lattice distance from the physical qubits on the boundary. After some number of known (heralded) erasures of the physical qubits has occured, one starts to reconstruct the logical qubits closest to the boundary, applying the decoding map corresponding to each individual tensor (e.g., of the $\llbracket 5,1,3\rrbracket$ code for the perfect-tensor pentagon code). Once these are (partially) reconstructed, one has access to the “intermediate” physical qubits corresponding to contracted legs one layer deeper into the bulk, repeating the local reconstruction procedure there then repeating the procedure. When no more logical qubits can be reconstructed, the greedy algorithm terminates and one has effectively found a decoder for logical qubits in a greedy wedge of the bulk on the non-erased physical qubits.
A greedy decoder can be straightforwardly constructed for any tensor network stabilizer code by translating possible operator pushing steps, which follow from the application of the stabilizers and gauge-fixed logical operators, into greedy steps. For any Evenbly code, for example, the operator pushing steps of Figure 4 (which are equivalent to the generalized isometry conditions of Figure 5) imply a greedy recovery for single and bulk-neighboring logical qubits that is shown in Figure 12(a)-(b). Note that this generalizes the original greedy algorithm of [8], in which each greedy step only recovers one logical qubit at a time.
Beyond these general recovery step that any Evenbly code fulfills, the GHZ-based code discussed in the main text allows for an additional greedy step shown in Figure 12(c), whereby access to physical sites on a tensor can assist in the recovery of the logical qubit on its neighbor.
In spite of these limiting factors, greedy recovery for a max-rate Evenbly code is generally much weaker than for a HaPPY code, owing to the effect that low-weight errors can have deep in the bulk [24], precluding the HaPPY codes’ feature of “uberholography” [40]. A generic example is given in Figure 12(d), where the erasure of only a few boundary sites prevents greedy reconstruction of most of the bulk.
@@ E.2 Gaussian Elimination Algorithm
After obtaining the binary reduced-row echelon form (RREF) $BSV^{\prime}_{\text{RREF}}$, we iterate through the rows of $BSV^{\prime}_{\text{RREF}}$ and check the last two columns. If either of the entries are of value 1, then we check the other values in the other (n2)(n-2) columns. If any value is equal to 1, then the algorithm succeeds, and we return $\mathrm{True}$ after incrementing the counter num_true. The algorithm terminates when all trials have completed, and the number of $\mathrm{True}$ values are counted against the number of trials (which we name $\mathcal{P}_{\text{rec}}=\frac{\text{num\_true}}{\text{mc\_trials}}$). This mathematical procedure largely follows [9].
Layer HTN 1 Extra Layer 2 Extra Layers
L = 0 $1.0e5$ $1.0e5$ $1.0e5$
L = 1 $1.0e5$ $1.0e5$ $1.0e5$
L = 2 $1.0e4$ $1.0e4$ $1.0e4$
L = 3 $1.0e3$ $1.0e3$ $1.0e3$
L = 4 $1.0e3$ $1.0e3$ $1.0e3$
@@
In the simulations executed for this work, we vary the number of Monte Carlo simulations as a function of the number of layers, in order to efficiently delegate computational resources. The number of Monte Carlo trials per layer and variant tested are shown in Table 2. The first column, labeled ”HTN”, corresponds to the zero-rate and constant-rate Evenbly codes tested in Sections 3.5 and 3.6, while the others refer to the other variants of the constant-rate HTN in which extra layers of Z or $Y$ gauge are added.
Input: $\mathcal{S}_{R},\mathcal{L}_{R},n,\text{mc\_trials}$
$\text{BSV}\leftarrow\text{convert\_to\_BSV}(\mathcal{S}_{R},\mathcal{L}_{R})$
$\text{num\_true}=0$
for $\text{trial}\in\text{mc\_trials}$ do
@@
return $\mathcal{P}_{\text{rec}}=\text{num\_true}/\text{mc\_trials}$